1. From a regression analysis of sales with respect to advertising expenses in a study for a department store chain, the following output is obtained (use 0.05 significance level):
SUMMARY OUTPUT
Regression Statistics
Multiple R 0.954
R Square 0.91
Adjusted R Square 0.902
Standard Error 936.85

ANOVA
df SS MS F
Regression 1 10621 10621 121.01
Residual 12 1053.2 87.768
Total 13 11674

Coefficients Standard Error t Stat P-value
Intercept 901.25 513.02 1.757 0.104
Advertising 1.686 0.153 11 0.000

(e) What is the regression equation?




(c) Is advertising expenses a significant factor for sales? Please give reason.




(f) Predict the sales volume if $5000 is spent on advertising.



(g) Determine the correlation coefficient



(h) Provide an interpretation of the correlation coefficient

Answer by stanbon(48568)

1. From a regression analysis of sales with respect to advertising expenses in a study for a department store chain, the following output is obtained (use 0.05 significance level):
SUMMARY OUTPUT
Regression Statistics
Multiple R 0.954
R Square 0.91
Adjusted R Square 0.902
Standard Error 936.85
ANOVA
df SS MS F
Regression 1 10621 10621 121.01
Residual 12 1053.2 87.768
Total 13 11674
Coefficients Standard Error t Stat P-value
Intercept 901.25 513.02 1.757 0.104
Advertising 1.686 0.153 11 0.000
(e) What is the regression equation?
Sales = 1.686(ad expenses) + 901.25
-----------------------------------------------
(c) Is advertising expenses a significant factor for sales? Please give reason.
Yes ; the p-value of 0.000 indicates that a hypothesis test of the correlation
of sales and add expenses would reject Ho if Ho denied a linear relationship.
------------------
(f) Predict the sales volume if $5000 is spent on advertising.
s(5000) = 1.688(5000)+ 901.25
-----------------
(g) Determine the correlation coefficient
r = sqrt(R^2) = sqrt(0.91) = 0.9539
--------------
(h) Provide an interpretation of the correlation coefficient
95.29% of the variation between sales volume and advertising
is explained in the regression equation.
=====================================================
Cheers,
Stan H.